### v1

# Fast, accurate and lightweight sequential simulation-based inference using Gaussian locally linear mappings

Abstract
Bayesian inference for complex models with an intractable likelihood can be tackled using algorithms performing many calls to computer simulators. These approaches are collectively known as simulation-based inference (SBI). Recent SBI methods have made use of neural networks (NN) to provide approximate, yet expressive constructs for the unavailable likelihood function and the posterior distribution. However, they do not generally achieve an optimal trade-off between accuracy and computational demand. In this work, we propose an alternative that provides both approximations to the likelihood and the posterior distribution, using structured mixtures of probability distributions. Our approach produces accurate posterior inference when compared to state-of-the-art NN-based SBI methods, while exhibiting a much smaller computational footprint. We illustrate our results on several benchmark models from the SBI literature.
${}^{1}$ept. Mathematical Sciences, Chalmers and University of Gothenburg ${}^{2}$niv. Grenoble Alpes, Inria, CNRS, Grenoble INP, Inserm U1216, CHU Grenoble Alpes, Grenoble Institut des Neurosciences, LJK ${}^{3}$niv. Lyon, CNRS, Inserm, INSA Lyon, UCBL, CREATIS Correspondence to: picchini@chalmers.se
1 Introduction
We propose a novel methodology achieving state-of-the-art Bayesian inference for models with an intractable likelihood function. We illustrate, via several benchmark models, how our methodology returns accurate inference, even in presence of multimodal posteriors, with a small runtime. Moreover, for numerous simulation studies and applications, machine and statistical learning solutions show an ever growing performance, but are not generally evaluated in terms of resource and energy consumption. Although their environmental impact is more and more documented (Schwartz et al., 2020; Strubell et al., 2019; Thompson et al., 2022), the question of resource efficiency and energy consumption when learning model parameters is not as much studied as the quality of the inference itself, due to the multiple aspects to be taken into account and to the lack of clear metrics that limit the algorithms evaluation. When comparing with widely employed inference engines based on neural-networks, our method considerably lower the energy and memory required for the inference task.
We consider simulation-based inference (SBI) for model parameters $\bm{\theta}$ in stochastic modelling. SBI refers to approaches that bypass the need to have a likelihood function $p(\bm{y}|\bm{\theta})$ (for data $\bm{y}$) available in closed form, or one that is cheap to approximate, by instead considering a generative model $f(\bm{\theta})$ or computer simulator. The simulator depends on $\bm{\theta}$ and possibly on other inputs such as covariates and a stream of pseudorandom numbers. Generating a dataset $\bm{y}^{*}$ by running the simulator $\bm{y}^{*}=f(\bm{\theta}^{*})$ at $\bm{\theta}=\bm{\theta}^{*}$ is then equivalent to simulating $\bm{y}^{*}$ from the unavailable likelihood function. In the following this simulation is written as $\bm{y}^{*}\sim p(\bm{y}|\bm{\theta}^{*})$. Established SBI methods include approximate Bayesian computation (ABC, reviewed in Sisson et al., 2018), synthetic likelihoods (Wood, 2010; Price et al., 2018), particle Markov chain Monte Carlo (Andrieu and Roberts, 2009; Andrieu et al., 2010), and several more that we discuss in Section 2. More recently, methods exploiting neural networks (NN) have gained considerable attention. Neural-based estimators have been used to produce approximated likelihoods or posterior distributions, typically using (discrete) normalizing flows (Rezende and Mohamed, 2015; Papamakarios et al., 2021). NN have been used to approximate the likelihood function in implicit models (Papamakarios et al., 2019; Chen et al., 2021), the posterior distribution (Papamakarios and Murray, 2016; Greenberg et al., 2019; Durkan et al., 2020; Chen et al., 2021) or the likelihood and the posterior distribution simultaneously (Wiqvist et al., 2021; Radev et al., 2023a).
These neural-based posterior and likelihood estimation methods have produced exceptionally flexible inference strategies for models with intractable likelihoods. In this work, we propose to investigate more frugal strategies (Evchenko et al., 2021) that can run with limited resources and a much smaller computational footprint, while returning inference of similar quality as neural-based approaches. For SBI, we propose to investigate structured mixtures of probability distributions whose versatility has been widely recognized for a variety of data and tasks, while not requiring excessive design effort or tuning. Their expressivity makes them good candidates to account for complex multivariate models both for the likelihood and posterior distribution. Moreover, mixture models have a much smaller number of hyperparameters compared to NN, and this makes them more amenable to interpretation and efficient learning (see Table 1). More specifically, we design a new simulation-based inference method named Sequential Mixture Posterior and Likelihood Estimation (SeMPLE). SeMPLE learns approximations for both the posterior $p(\bm{\theta}|\bm{y})$ and the likelihood $p(\bm{y}|\bm{\theta})$ using a structured mixture model obtained via the Gaussian Locally Linear Mapping (GLLiM, Deleforge et al., 2014) method. Interestingly, both approximations are jointly returned following a run of an Expectation-Maximization (EM) algorithm within GLLiM. We then use the GLLiM approximate posterior as a well-informed proposal distribution in a tuning-free Metropolis-Hastings sampler. The procedure is embedded in a sequential strategy, which we compare with state-of-the-art NN-based sequential algorithms for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Other algorithms could be considered for comparison, however we stick with SNL and SNPE-C and a motivation for this is given in Section 2. Our comparisons are in terms of inference quality as well as resource and energy requirements. In general, we did not compare against ABC samplers such as SMC-ABC, except for the Lotka-Volterra example where such comparison was more suitable, as in Lueckmann et al., 2021 it was shown that standard implementations of ABC approaches usually require a larger number of model simulations, compared to NN-based approaches, although recent research improved the SMC-ABC performance (Picchini and Tamborrino, 2022). We found that, for a given computational budget specified by the total number of model simulations, our method is cheaper to run and produces accurate inference thanks to (i) the informed proposal sampler created by GLLiM and (ii) the computationally faster model training. As an illustration, Figure 1 shows inference for three datasets of the Two Moons model, a model with bimodal posterior that we further analyse in Section 5. The runtime, energy and memory requirements for SeMPLE are considerably lower, see Table 1 and Figure 2. Overall SeMPLE shows low resources costs that are always of the same magnitude ($<10$ kJ and $<1$ GB), while the requirements for SNL and SNPE-C vary much more with the model dimension or the inference complexity. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance.
Figure 1: Two Moons model from SBIBM, datasets 1 (top), 8 (middle) and 9 (bottom). Posterior samples after 4 rounds of SeMPLE (column 1), 10 rounds of SNPE-C and SNL (columns 2 and 3), using 10K model simulations. The true posteriors are in column 4.
Table 1: Median resource requirements (across 10 independent repetitions) and number of model hyperparameters $\phi$ for posterior inference with a budget of 10K (two moons and Bernoulli GLM) and 40K (hyperboloid and Ornstein-Uhlenbeck) model simulations. Best (lowest) values are in bold. SNPE-C and SNL used the default SBIBM setup.
Method CPU+DRAM Peak Memory Usage #Hyperparameters
Two moons
SNL 24.5 kJ 19 GB 28,020
SNPE-C 15 kJ 9 GB 60,925
SeMPLE 7.1 kJ 0.71 GB 449
Hyperboloid
SNL 54 kJ 30 GB 30,020
SNPE-C 120 kJ 15 GB 66,925
SeMPLE 3.2 kJ 0.88 GB 1,479
Bernoulli GLM
SNL 108 kJ 75 GB 36,100
SNPE-C 41 kJ 10 GB 98,025
SeMPLE 4.8 kJ 0.93 GB 2,309
Ornstein-Uhlenbeck
SNL 66 kJ 39 GB 41,030
SNPE-C 134 kJ 15 GB 102,242
SeMPLE 6.1 kJ 0.93 GB 30,799
Section 2 gives a brief overview of the main SBI approaches. Section 3 describes how GLLiM can be used to construct surrogate posteriors and surrogate likelihoods, while the complete SeMPLE methodology is explained in Section 4. Results from SeMPLE, SNL and SNPE-C are then illustrated in Section 5. Test models, partly reported in the Supplementary Material, include six simulation studies comprising multimodal surfaces (two moons, multiple hyperboloid), stochastic differential equations (Ornstein-Uhlenbeck), generalised linear models (Bernoulli GLM), a Markov jump process with additional measurement error (Lotka-Volterra), and highly correlated posteriors with 20 parameters (twisted-prior). Code is available at https://github.com/henhagg/semple.
2 Related work
Simulation-based inference, also referred to as likelihood-free inference, is reviewed e.g. in Cranmer et al. (2020) and Pesonen et al. (2023). In a Bayesian framework the focus is on posterior estimation, which in SBI directly relates to the quality of the approximations made to the unavailable likelihood, the posterior distribution or both. This aspect raises the issue of how to evaluate the quality of posterior inference when a reference ground truth posterior is absent by definition. This is still an open question and is further discussed in Section 6. A second important ingredient is sampling efficiency, as simulation-based methods are typically relying on exploration, and on Monte Carlo principles that can be prohibitively expensive. Active or sequential learning is a way to address sample efficiency by gradually discovering and guiding where simulations should be made for a specific given observation. Our proposed method is sequential and we focus on comparisons with NN-based sequential methods for posterior inference, namely SNL (Papamakarios et al., 2019) and SNPE-C (Greenberg et al., 2019). Many other approaches exist for Bayesian inference with intractable likelihoods, such as sequential neural ratio estimation (SNRE, Hermans et al., 2020), the several variants of ABC such as SMC-ABC (Toni et al., 2009; Del Moral et al., 2012), synthetic likelihoods (Wood, 2010; Price et al., 2018) and particle MCMC (Andrieu and Roberts, 2009; Andrieu et al., 2010). Comparisons between several of the above mentioned methods are available in previous works. For example, Greenberg et al. (2019) show that SNPE-C is computationally more efficient than the predecessors [^1] The acronyms SNPE-A and SNPE-B follow the taxonomy in Papamakarios et al. (2019). SNPE-A (Papamakarios and Murray, 2016) and SNPE-B (Lueckmann et al., 2017), where efficient means that, for a given computational budget, SNPE-C approaches the true posterior using a smaller number of model simulations. Lueckmann et al. (2021) show that, for nontrivial case-studies, both SNPE-C and SNRE are more efficient than SNL and SMC-ABC. Moreover, SNPE-C and SNRE perform similarly. The above motivates why in this work we focus on the comparison with SNPE-C using the SBIBM package implementation (Lueckmann et al., 2021) (an exception is the Lotka-Volterra Markov jump process, where we compare against SMC-ABC, as an objective comparison with NN-based methods is difficult here, see section 5.3 for more details). Since SeMPLE includes an MCMC step, same as SNL, it is also relevant to compare against SNL, using its implementation in SBIBM, which is an ameliorated one, compared to the original version of Papamakarios et al. (2019).
Note that there are similarities between our SeMPLE and SNPE-A, as detailed in Supplementary Section 3. Briefly, in SNPE-A a $K$ component mixture density network (MDN) is used as a proposal sampler. As such, the MDN means and covariances are parametrised via neural networks. However SNPE-A does not have a built-in strategy to learn each of the $K$ components. Instead, the GLLiM procedure embedded into SeMPLE automatically determines the means and covariance matrices for each of the $K$ components, thus increasing flexibility to the shapes of the posteriors that can be targeted. The fact that SeMPLE does not use MDNs is key to its resource-efficiency, since no complex NN architectures are employed in SeMPLE, and the benefit of the EM algorithm can be exploited towards fitting the mixture models.
Recent work that is relevant for SBI, but is less comparable with our work, includes amortized strategies. Amortization is a property that refers to the ability of algorithms to make use of previous computational efforts, not to repeat computationally expensive steps for each new observation considered. Examples in this class are the approaches in Radev et al. (2023a), Radev et al. (2023b), Dyer et al. (2022), Gao et al. (2023). Although GLLiM is by design an amortized procedure, this property is partly lost when used in a sequential approach such as SeMPLE, as amortization and sequential learning are somewhat opposite objectives. The next section gives a brief introduction to the Gaussian local linear mapping (GLLiM) procedure, which allows to obtain amortized surrogates of the posterior and of the likelihood in closed form.
3 Amortized surrogates for likelihoods and posteriors via GLLiM
The inverse regression approach via Gaussian local linear mapping (GLLiM) was first introduced in Deleforge et al. (2014), and has been employed for inference via approximate Bayesian computation (ABC) in Forbes et al. (2022). GLLiM is a parametric mixture model on the joint distribution of $(\bm{y},\bm{\theta})$, denoted by $q_{\tilde{\bm{\phi}}}$ and depending on a set of parameters $\tilde{\bm{\phi}}$. For interpretation or visualization, the model can also be presented as an invertible regression model as follows. In GLLiM the relationship between observations $\bm{y}\in\mathbb{R}^{d}$ and $\bm{\theta}\in\mathbb{R}^{l}$ is assumed to be locally-linear with $K$ components, defined through an additional latent variable $z\in\{1,...,K\}$ as,
$\bm{y}=\sum_{k=1}^{K}\mathbbm{1}\{z=k\}(\tilde{\bm{A}}_{k}\bm{\theta}+\tilde{\bm{b}}_{k}+\tilde{\bm{\epsilon}}_{k}),$ (1)
where $\mathbbm{1}$ denotes the indicator function, and $\tilde{\bm{A}}_{k}\in\mathbb{R}^{d\times l}$ and $\tilde{\bm{b}}_{k}\in\mathbb{R}^{d}$ are matrices and vectors, respectively, defining the affine transformations, while $\tilde{\bm{\epsilon}}_{k}\in\mathbb{R}^{d}$ corresponds to an error term capturing both the observational noise and the modelling error due to assuming an affine approximation for the data. In the following we consider a Gaussian noise, $\tilde{\bm{\epsilon}}_{k}\sim\mathcal{N}_{d}(\bm{0},\tilde{\bm{\Sigma}}_{k}$), and assume $\tilde{\bm{\epsilon}}_{k}$ not depending on $\bm{\theta}$, $\bm{y}$ nor $z$. Therefore, an approximate generative model (while the assumed generative model is denoted $p(\bm{y}|\bm{\theta})$) is given by
$q_{\tilde{\bm{\phi}}}(\bm{y}\,|\,\bm{\theta},z=k)=\mathcal{N}_{d}(\bm{y};\tilde{\bm{A}}_{k}\bm{\theta}+\tilde{\bm{b}}_{k},\tilde{\bm{\Sigma}}_{k}).$ (2)
To complete the hierarchical model, $\bm{\theta}$ is assumed to follow a mixture of Gaussian distributions specified by
$q_{\tilde{\bm{\phi}}}(\bm{\theta}\,|\,z=k)$ $=\mathcal{N}_{l}(\bm{\theta};\tilde{\bm{c}}_{k},\tilde{\bm{\Gamma}}_{k}),$ (3a)
$q_{\tilde{\bm{\phi}}}(z=k)$ $=\pi_{k}.$ (3b)
The GLLiM model hierarchical construction above (eq.(1) to (3b)) defines then a joint Gaussian mixture model on $(\bm{y},\bm{\theta})$:
$q_{\tilde{\bm{\phi}}}(\bm{y},\bm{\theta})\!=\!\sum_{k=1}^{K}q_{\tilde{\bm{\phi}}}(\bm{y}\!\,|\,\!\bm{\theta},z\!=\!k)q_{\tilde{\bm{\phi}}}(\bm{\theta}\!\,|\,\!z\!=\!k)q_{\tilde{\bm{\phi}}}(z\!=\!k),$ (4)
where the full vector of mixture model parameters is
$\tilde{\bm{\phi}}=\{\pi_{k},\tilde{c}_{k},\tilde{\bm{\Gamma}}_{k},\tilde{\bm{A}}_{k},\tilde{\bm{b}}_{k},\tilde{\bm{\Sigma}}_{k}\}_{k=1}^{K}.$ (5)
From this joint model, we can deduce straightforwardly both its conditional distributions $q_{\tilde{\bm{\phi}}}(\bm{y}\,|\,\bm{\theta})$ and $q_{{\bm{\phi}}}(\bm{\theta}\,|\,\bm{y})$ in closed-form. First, it comes that,
$q_{\tilde{\bm{\phi}}}(\bm{y}\,|\,\bm{\theta})=\sum_{k=1}^{K}\tilde{\eta}_{k}(\bm{\theta})\mathcal{N}_{d}(\bm{y};\tilde{\bm{A_{k}}}\bm{\theta}+\tilde{\bm{b}}_{k},\tilde{\bm{\Sigma}}_{k}),$ (6)
with
$\tilde{\eta}_{k}(\bm{\theta})=\frac{\pi_{k}\mathcal{N}_{l}(\bm{\theta};\tilde{\bm{c}}_{k},\tilde{\bm{\Gamma}}_{k})}{\sum_{j=1}^{K}\pi_{j}\mathcal{N}_{l}(\bm{\theta};\tilde{\bm{c}}_{j},\tilde{\bm{\Gamma}}_{j})},$ (7)
and similarly,
$q_{{\bm{\phi}}}(\bm{\theta}\,|\,\bm{y})=\sum_{k=1}^{K}\eta_{k}(\bm{y})\mathcal{N}_{l}(\bm{\theta};\bm{A}_{k}\bm{y}+\bm{b}_{k},\bm{\Sigma}_{k}),$ (8)
with
$\eta_{k}(\bm{y})=\frac{\pi_{k}\mathcal{N}_{d}(\bm{y};\bm{c}_{k},\bm{\Gamma}_{k})}{\sum_{j=1}^{K}\pi_{j}\mathcal{N}_{d}(\bm{y};\bm{c}_{j},\bm{\Gamma}_{j})},$ (9)
where $\bm{\phi}=\{\pi_{k},\bm{c}_{k},\bm{\Gamma}_{k},\bm{A}_{k},\bm{b}_{k},\bm{\Sigma}_{k}\}_{k=1}^{K}$ is just a convenient equivalent reparameterization that can easily be deduced from $\tilde{\bm{\phi}}$ (Deleforge et al., 2014) as
$\bm{c}_{k}$ $=\tilde{\bm{A}_{k}}\tilde{\bm{c}}_{k}+\tilde{\bm{b}}_{k}$ (10a)
$\bm{\Gamma}_{k}$ $=\tilde{\bm{\Sigma}}_{k}+\tilde{\bm{A}_{k}}\tilde{\bm{\Gamma}}_{k}\tilde{\bm{A}_{k}}^{\top}$ (10b)
$\bm{\Sigma}_{k}$ $=\big{(}\tilde{\bm{\Gamma}}_{k}^{-1}+\tilde{\bm{A}_{k}}^{\top}\tilde{\bm{\Sigma}}_{k}^{-1}\tilde{\bm{A}_{k}}\big{)}^{-1}$ (10c)
$\bm{A}_{k}$ $=\bm{\Sigma}_{k}\tilde{\bm{A}_{k}}^{\top}\tilde{\bm{\Sigma}}_{k}^{-1}$ (10d)
$\bm{b}_{k}$ $=\bm{\Sigma}_{k}\big{(}\tilde{\bm{\Gamma}}_{k}^{-1}\tilde{\bm{c}}_{k}-\tilde{\bm{A}_{k}}^{\top}\tilde{\bm{\Sigma}}_{k}^{-1}\tilde{\bm{b}}_{k}\big{)}.$ (10e)
The modelling power of GLLiM comes essentially from the fact that both conditional distributions above are mixtures of Gaussian affine experts (MoE), which are known for their modelling and approximation capabilities (Nguyen et al., 2019, 2021, 2023). As such, they are good candidates to provide surrogate likelihoods and posteriors. All the more so as, given a training dataset made of many pairs of $\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$, surrogate models specified via GLLiM can be estimated via an EM algorithm, see Deleforge et al. (2014) and Forbes et al. (2022) for details. In fact a surrogate likelihood and posterior can be obtained with $\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$ prior-predictive realizations simulated from $p(\bm{\theta})p(\bm{y}\,|\,\bm{\theta})$, but more generally, as exploited in SeMPLE, GLLiM can be used in a sequential way, by changing the training set to include posterior draws.
Once an estimate for the parameter vector $\tilde{\bm{\phi}}$ (or equivalently $\bm{\phi}$) is available, we have an amortized procedure providing a surrogate likelihood function $q_{\tilde{\bm{\phi}}}(\bm{y}\,|\,\bm{\theta})$, which can be evaluated for any data $\bm{y}$. In particular, if we specify $\bm{y}$ to be a given observed dataset $\bm{y}_{o}$, and plug $\bm{y}_{o}$ into (6), we obtain an approximate (surrogate) likelihood function $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$ for our observations $\bm{y}_{o}$. Similarly, an amortized surrogate posterior, $q_{\bm{\phi}}(\bm{\theta}\,|\,\bm{y})$, for a generic $\bm{y}$, can also be derived as a mixture of Gaussian experts distributions from (8). If we plug observed data into (8), then $q_{\bm{\phi}}(\bm{\theta}\,|\,\bm{y}_{o})$ is an approximate posterior based on data. In summary, once $\tilde{\bm{\phi}}$ is estimated via EM, we obtain surrogates of both the likelihood and the posterior in the same GLLiM run. The way this is performed and integrated with the rest of the SeMPLE method is explained in Section 4. Note that it is possible to assume a different distribution for $\tilde{\bm{\epsilon}}_{k}$ than the Gaussian one. For example (1) can be a mixture of generalized Student distributions, resulting in the so-called SLLiM method (Perthame et al., 2018). Both GLLiM and SLLiM are implemented in the R package xLLiM (Perthame et al., 2022) which we employ in our experiments, see Section 5 for details.
Algorithm 1 SeMPLE
1:$\tilde{p}_{0}(\bm{\theta})\leftarrow p(\bm{\theta})$, set starting $K$ and number of rounds $R$.
2:for $r=0:R$ do
3:     for $n=1:N$ do
4:         if $r=0$ or $r=1$ sample $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ without MCMC, otherwise sample $\bm{\theta}_{n}\sim\tilde{p}_{r}(\bm{\theta})$ using Algorithm 2.
5:         Sample (implicitly) $\bm{y}_{n}\sim p(\bm{y}\,|\,\bm{\theta}_{n})$.
6:     end for
7:     if r = 0 then
8:         Collect $\mathcal{D}_{0}=\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}$.
9:         Train $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{0}$. Update $K$.
10:         Obtain $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$.
11:         Set $\tilde{p}_{1}(\bm{\theta})\leftarrow q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}=\bm{y}_{o})$.
12:     else
13:         Collect $\mathcal{D}_{r}=\mathcal{D}_{r-1}\cup\{\bm{\theta}_{n},\bm{y}_{n}\}_{n=1}^{N}\setminus\mathcal{D}_{0}$.
14:         Train $q_{\tilde{\bm{\phi}}_{r}}(\bm{y}\,|\,\bm{\theta})$ on $\mathcal{D}_{r}$. Update $K$.
15:         Obtain $q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y})$.
16:         Set $\tilde{p}_{r+1}(\bm{\theta})\leftarrow q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$  ((optional) $\tilde{p}_{r+1}(\bm{\theta})\leftarrow\frac{p(\bm{\theta})}{\tilde{p}_{r}(\bm{\theta})}q_{\bm{\phi}_{r}}(\bm{\theta}\,|\,\bm{y}_{o})$)
17:     end if
18:end for
Algorithm 2 Independence MH with GLLiM surrogate posterior as proposal distribution
1:For the initial value $\bm{\theta}_{1}$, pick the last accepted value from the currently available Markov chain.
2:for $j=2:N$ do
3:     Propose $\bm{\theta}^{*}\sim q_{\bm{\phi}_{r-1}}(\bm{\theta}\,|\,\bm{y}_{o})$,
4:     $\alpha=\min\{1,\frac{\tilde{p}_{r}(\bm{\theta}^{*})}{\tilde{p}_{r}(\bm{\theta}_{j-1})}\frac{q_{\bm{\phi}_{r-1}}(\bm{\theta}_{j-1}\,|\,\bm{y}_{o})}{q_{\bm{\phi}_{r-1}}(\bm{\theta}^{*}\,|\,\bm{y}_{o})}\}.$
5:     Sample $u\sim\mathcal{U}[0,1]$
6:     if $u\leq\alpha$ then
7:         $\bm{\theta}_{j}=\bm{\theta}^{*}$,
8:     else
9:         $\bm{\theta}_{j}=\bm{\theta}_{j-1}$.
10:     end if
11:end for
4 SeMPLE: sequential learning of likelihoods and posteriors
We have shown how GLLiM can provide amortized estimations of both the likelihood and the posterior distribution at no additional cost, given training data. This GLLiM feature is exploited in our proposed SeMPLE procedure, to obtain a sequence of increasingly improved samplers, to more accurately and more efficiently sample from the targeted posterior of observed data $\bm{y}_{o}$. Similarly to SNL (Papamakarios et al., 2019), given a likelihood approximation $q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})$, SeMPLE targets an approximate posterior $\tilde{p}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ with posterior sampling being carried out using a Markov Chain Monte Carlo (MCMC) procedure. Typically, the MCMC scheme used in SNL is slice sampling (Neal, 2003), which has the advantage to be tuning-free but is likely to have difficulties in exploring high dimensional posteriors with complex multimodalities. As a matter of fact, Papamakarios et al. (2019) recommend SNL “at most for tens of parameters that do not have pathologically strong correlations”. In SeMPLE, since GLLiM is able to learn a closed-form approximation of the posterior distribution (in addition to a likelihood function), such posterior is used as a proposal distribution within a Metropolis-Hastings (MH) algorithm. SeMPLE, by using an efficient EM procedure to fit a flexible proposal distribution, is expected to benefit from a MH step that better guides simulations towards more likely parameter values. To ease the comparison with SNL, the latter is recalled in Supplementary Section 4, while SeMPLE is summarized in Algorithms 1-2.
Algorithm 1 shows that the targeted approximate posterior for $\bm{\theta}$ is a function denoted with $\hat{p}_{r}(\bm{\theta})\propto q_{\tilde{\bm{\phi}}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})p(\bm{\theta})$ which is sequentially obtained, where $r=0,...,R$ denotes the SeMPLE rounds for the current approximation [^2] We use the term round instead of the most common iteration, as the latter could be confused with the MCMC iterations that are run within a given SeMPLE round.. In the following we denote by $\bm{\theta}^{(r)}$ and $\bm{y}^{(r)}$ a parameter and an observation simulated in round $r$, respectively. At first ($r=0$) $\hat{p}_{0}(\bm{\theta})$ is initialised at the prior $p(\bm{\theta})$. An initial batch of $N$ pairs $\mathcal{D}_{0}=\{\bm{\theta}_{n}^{(0)},\bm{y}_{n}^{(0)}\}_{n=1}^{N}$ produced from the prior-predictive distribution (steps 4-5) constitutes the starting training data for GLLiM, and the parameters $\tilde{\phi}_{0}$ and ${\phi}_{0}$ are obtained via EM within GLLiM to produce the initial surrogate likelihood and posterior $q_{\tilde{\bm{\phi}}_{0}}(\bm{y}\,|\,\bm{\theta})$ and $q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y})$ respectively (steps 9-10). The initial surrogate posterior becomes the current target (step 11) $\tilde{p}_{1}(\bm{\theta})\equiv q_{\bm{\phi}_{0}}(\bm{\theta}\,|\,\bm{y}_{o})$. In the next iteration ($r=1$), the latter can be sampled from without MH (step 4) since $\tilde{p}_{1}(\bm{\theta})$ is defined as a Gaussian mixture, from which it is trivial to sample $N$ parameters $\{\bm{\theta}_{n}^{(1)}\}_{n=1}^{N}$. These $N$ draws are then plugged into the model simulator to produce $\bm{y}_{n}^{(1)}\sim p(\bm{y}\,|\,\bm{\theta}_{n}^{(1)})$ and define $\mathcal{D}_{1}=\{\bm{\theta}_{n}^{(1)},y_{n}^{(1)}\}_{n=1}^{N}$ as the new training data (we discard $\mathcal{D}_{0}$ for future iterations as it is deemed too uninformative for training). GLLiM is then fitted to $\mathcal{D}_{1}$ and the corresponding surrogate likelihood and surrogate posterior are obtained. From this step on, the surrogate posterior is employed as a proposal distribution within a tuning-free MH algorithm (Algorithm 2) since, for an arbitrary prior $p(\bm{\theta})$, when $r>1$ we do not explicitly know which distribution is proportional to the product $p(\bm{\theta})q_{\tilde{\phi}_{r}}(\bm{y}_{o}\,|\,\bm{\theta})$ (for $r=1$ we know by construction that the targeted posterior is a Gaussian mixture), except for the special case of Gaussian and Uniform priors (see Papamakarios and Murray, 2016). In Algorithm 2, the acceptance probability of a proposed $\bm{\theta}^{*}$ is (for $r>1$)
$\alpha=\min\biggl{\{}1,\frac{\tilde{p}_{r}(\bm{\theta}^{*})}{\tilde{p}_{r}(\bm{\theta}_{j-1})}\frac{q_{\bm{\phi}_{r-1}}(\bm{\theta}_{j-1}\,|\,\bm{y}_{o})}{q_{\bm{\phi}_{r-1}}(\bm{\theta}^{*}\,|\,\bm{y}_{o})}\biggr{\}}.$
Correlated samples from the posterior $\tilde{p}_{r}$ are obtained by first proposing parameters $\bm{\theta}^{*}$ from a Gaussian mixture model. From the current proposal function $q_{\phi_{r-1}}(\bm{\theta}\,|\,\bm{y}_{o})$ which is the mixture
$\sum_{k=1}^{K}\eta_{k}^{(r-1)}(\bm{y}_{o})\mathcal{N}_{l}(\bm{\theta};\bm{A}_{k}^{(r)}\bm{y}_{o}~{}\!\!+\!\!~{}\bm{b}_{k}^{(r-1)},\bm{\Sigma}_{k}^{(r-1)}),$
a component $k^{*}$ is randomly sampled from the set $\{1,...,K\}$ with associated probabilities $\{\eta_{1}^{(r-1)}(\bm{y}_{0}),...,\eta_{K}^{(r-1)}(\bm{y}_{0})\}$. Then $\bm{\theta}^{*}$ is sampled from the selected component $\mathcal{N}_{l}(\bm{\theta};\bm{A}_{k^{*}}^{(r)}\bm{y}+\bm{b}_{k^{*}}^{(r)},\gamma\bm{\Sigma}_{k^{*}}^{(r)})$ by (optionally) slightly inflating its covariance matrix with a factor $\gamma\geq 1$, which we set to $\gamma=1$ (no inflation) or $\gamma\in[1.1,1.2]$ in our experiments. The inflation factor is to ensure that the proposal is sampled from a broad-enough distribution covering the support of the target (no inflation factor is used when $r=1$). Also, note that the $N$ retained draws via MH are $N$ post-burnin draws, where the burnin consists of 100 iterations, and the last accepted draw from round $r$ of SeMPLE is used to initialize the chain at round $r+1$.
Importantly, our MCMC procedure is tuning-free. We use an independence sampler (Robert and Casella, 2004), relying on GLLiM to provide an informative proposal function. This proposal function is trained on draws accepted in previous rounds which are, reasonably, more spread than the currently targeted posterior. In addition, being a mixture model, it is suitable to deal with multimodalities in the target. This greatly simplifies the construction of SeMPLE as no complex tuning is necessary for the MCMC sampler. While $K$ could in principle be set arbitrarily large, in practice this may slow the EM algorithm convergence. In SeMPLE, the value of $K$ is allowed to decrease between iterations in the update $K$ step lines 9 and 14 in Algorithm 1. Mixture components having probabilities $\tilde{\eta}_{k}$ smaller than a threshold are discarded. In most cases we find that setting a starting value of $K=20$ or 30 produces good results, with a threshold for $\tilde{\eta}_{k}$ typically set to 0.005 (multiple hyperboloid, Ornstein-Uhlenbeck, Lotka-Volterra models), 0.03 (Bernoulli GLM model) or simply 0 for no deletion (Two Moons model). This is to prevent having too many irrelevant components, that could be problematic when fitting the mixture to training data produced via MCMC, as the latter may not manage to visit tiny modes. It is also possible to choose the starting value of $K$ in a principled way. In Deleforge et al. (2014), the Bayesian Information Criterion (BIC) is used to select $K$, as we do in our examples, see Supplementary Section 2. An appropriate value for $K$ can be learned by repeatedly computing the BIC on realizations of the prior-predictive distribution, before SeMPLE is started. This is an amortized procedure, as the selected values can be reused regardless the specific observed data $\bm{y}_{o}$. The impact on the overall computational budget is negligible. We have also experimented with a version of SeMPLE that learns a corrected posterior (see the optional part in line 16 of Algorithm 1) using the correction formula from Papamakarios and Murray (2016). However, we found that learning a surrogate likelihood was much more beneficial, providing better inference coupled with a higher acceptance rate. Notice that none of our reported results use the optional corrected-posterior approach.
In summary, a major difference with the several variants of sequential neural posterior estimation (where the most used one is SNPE-C, Greenberg et al., 2019), and with SNL, is that SeMPLE uses Gaussian mixtures learned via EM to approximate both the likelihood function and the posterior distribution of $\bm{\theta}$, as opposed to neural density estimation. Thus no NN architecture has to be designed, nor special stochastic gradient descent (SGD) setup has to be considered to fit NNs. Only the number of components $K$ and the structure of the covariance matrices $\tilde{\bm{\Sigma}}_{k}$ (as the $\bm{\Sigma}_{k}$ are obtained as an algebraic by-product), according to the options offered by the xLLiM package [^3] The $\tilde{\bm{\Sigma}}_{k}$’s can be isotropic, diagonal or full matrices, and set all equal or varying with $k$., have to be set. As we show in our examples, inference based on procedures using SGD can be much more computationally demanding and time consuming.
5 Numerical illustrations
We illustrate SeMPLE on several simulation studies, see also the Supplementary Material for extra results and additional models. SeMPLE is written in R and uses the GLLiM implementation from the xLLiM package (Perthame et al., 2022). Comparisons with neural-based likelihood and posterior estimation methods (SNL and SNPE-C) use the Python implementations in the SBIBM package for SBI benchmarking (Lueckmann et al., 2021), and architecture specifications for SBIBM are in Supplementary Section 1. To compare approximate posteriors with a suitable reference posterior, we consider the classifier two-samples test (C2ST), as implemented in SBIBM, and Wasserstein distances via the POT Python package (Flamary et al., 2021). C2ST varies in the interval [0.5,1], the lower the better. It equals 0.5 when samples from two distributions appear statistically indistinguishable and equals 1 when the classifier is able to perfectly separate samples as belonging to two different distributions. We also report runtimes in figures and resource requirements in Table 1. For SeMPLE, the time required to select $K$ via BIC is reported separately and not taken into account in the figures below. For some case studies, such as Lotka-Volterra, we merely picked a starting value of $K$ without running the BIC procedure. Adding this overhead time does not change the overall conclusion that SeMPLE is much more time-efficient than the other methods. Moreover, the selection of $K$ is an amortized procedure, hence it does not need to be repeated for different observations of the same model.
In the next two sections 5.1 and 5.2, we illustrate results from two models having multimodal posteriors, in section 5.3 we summarise the Lotka-Volterra experiment, and in section 5.5 and Table 2 we summarise results from all other experiments detailed in Supplementary Material. For comparison, all methods are given the same budget of model simulations. For SNL and SNPE-C the SBIBM default of distributing these simulations uniformly across $R=10$ rounds is used, but to ease comparison we also run SNL and SNPE-C with the same number of rounds used for SeMPLE, which is $R=4$, except for the Bernoulli GLM model where $R=2$ suffices. In fact, we observe that SeMPLE does not need as many sequential steps as the other methods. It may be argued that SNL could perform better if a larger number of MCMC iterations was used, at the cost of a smaller number of rounds $R$. Results in Supplementary Material show that both SNL and SNPE-C perform worse under this alternative configuration. Regarding SeMPLE, when the computational budget limits the number of model simulations, we found it beneficial to keep the number of rounds $R$ between 2 and 3, where $R=2$ implies that we have exactly one round where MCMC is used, and this is typically already enough for satisfying inference, though usually an additional round with MCMC (hence $R=3$) produces further refined inference. In our work we show results up to $R=4$, but in general, four rounds are not recommended. The algorithm performance relies first on the GLLiM ability to learn reasonably accurate likelihood and posterior approximations and then on the MCMC sampler to explore the parameter space. Both steps require a large enough number of model simulations, which depends mainly on the model dimension.
5.1 Two Moons
The Two Moons model is a standard benchmark example in SBI (Greenberg et al., 2019; Lueckmann et al., 2021; Picchini and Tamborrino, 2022) to illustrate how algorithms deal with multimodality and the local crescent shapes of the posterior distribution. The model itself is defined in Supplementary Section 5. Our experiments are run on the same 10 datasets provided in SBIBM, each of them being a vector $\bm{y}_{0}$ of length two. We use $K=30$, a value we determine as described in Supplementary Material (this determination only requires 90 seconds). We run SeMPLE, SNL and SNPE-C algorithms for a total budget of $10^{4}$ model simulations, distributed across inference rouces, and we tested both $R=4$ and $R=10$ rounds. The obtained samples are then compared with the corresponding posterior reference samples (obtained via MCMC, since the likelihood is tractable) provided in SBIBM. The performance metric results are averaged over the 10 different data sets. Figure 2 shows C2ST and runtimes values. SeMPLE provides the best approximation while SNL performs the worst. The comparison with SNL is particularly interesting as while both SNL and SeMPLE incorporate a tuning-free MCMC step, SeMPLE is able to visit both modes with an acceptance rate of 60-70%, while SNL often struggles in efficiently exploring a multimodal posterior as demonstrated by the higher C2ST values. Moreover, in this simple example the runtime for SeMPLE is about half the SNPE-C runtime, and is about three times lower than the SNL runtime. In other examples (e.g. Figure 3) the speedup provided by SeMPLE is even more dramatic. All these benefits from SeMPLE come with a very small impact on the memory usage, see Section 5.4 for details.
Figure 2: Two Moons. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.
5.2 Multiple hyperboloid model
The multiple hyperboloid model (Forbes et al., 2022) is constructed from a sound source localization problem in audio processing. This is a challenging inference task with a complex symmetrical and multimodal posterior distribution. The model is provided in Supplementary Section 6. The observation is a vector of length 10 and $\bm{\theta}$ has dimension 2. SeMPLE is run with $K=40$, a value that is selected via BIC in 4.5 minutes. The total number of model simulations is set to $4\times 10^{4}$. C2ST and runtime results are reported in Figure 3. An exemplary posterior is shown in Figure 4. SeMPLE shows the best performance according to C2ST, and its runtime is about half that of SNL, and is about 10 times more efficient than SNPE-C. Adding the BIC overhead does not change the conclusion that SeMPLE is much more efficient. The runtime comparison with SNL is not particularly relevant, since the latter does not generally manage to explore the multimodal posterior (see Figure 4 and the posteriors in Supplementary Section 6).
Figure 3: Hyperboloid. Median C2ST (left) and median cumulative runtime in minutes (right) for 10 runs with different data sets vs the number of model simulations. Shaded bands enclose the min and max values.
Figure 4: Hyperboloid. An example of marginal posteriors and samples from the last iteration of each algorithm. The SNL marginal posteriors are not reported as their inference is completely off.
5.3 Lotka-Volterra
The Lotka-Volterra predator prey model is a classical example of a population dynamics model. We consider the stochastic Markov jump process version of the model, which is defined in Supplementary Section 9. This model, although being relatively simple and characterized by a set of three interactions on the two species, each with a corresponding rate $\theta_{j}$ ($j=1,2,3$), encompasses many of the difficulties associated with larger, more complex systems. The observations consider the addition of non-negligible measurement error to the simulated dynamics, and the standard deviation $\sigma$ of this noise term is inferred as an additional parameter, resulting in a total of four model parameters to infer $(\theta_{1},\theta_{2},\theta_{3},\sigma)$. For this model, we are not primarily focusing on the performance comparison to other neural-based methods, as this version of the model, as a Markov jump process, is not included in SBIBM. Moreover, running times for a single model call are random even for repeated calls using the same parameter value. We use the Gillespie algorithm as detailed in Supplementary Section 9, and the number of “reactions” at any given $\bm{\theta}$ is stochastic. Due to the reasons above, the primary focus is to show that high quality inference can be obtained with SeMPLE using a comparably small computational effort, and compare our results against two versions of sequential Monte Carlo ABC (SMC-ABC). A “reference posterior” is obtained with an efficient version of SMC-ABC called blockedopt in Picchini and Tamborrino (2022), which is run for very many model simulations, around 3.3 million, see the Supplementary Material for details. We also compare it with a more standard off-the-shelf SMC-ABC named standard. SeMPLE inference is obtained with only 30,000 model simulations. In Figure 5 we also include the corresponding inference obtained with around 30,000 model simulations for both blockedopt and standard SMC-ABC. With only 30,000 model simulations, SeMPLE manages to produce inference for all parameters similar to the inference returned after 3.3 million model simulations from blockedopt SMC-ABC, a 100-fold computational saving. We observe that, unlike SMC-ABC, SeMPLE produces accurate posterior inference for all model parameters, including the noise parameter $\sigma$, but with a much smaller number of model simulations compared to SMC-ABC. The latter fact, namely the SMC-ABC inefficiency in terms of model simulations, was not unexpected and known from Greenberg et al. (2019); Lueckmann et al. (2021). However it is reassuring to notice the excellent inference quality brought by SeMPLE in this challenging case study.
Figure 5: Lotka-Volterra. Marginal posteriors from SMC-ABC and SeMPLE. SeMPLE is run for a total of 30,000 model simulations. We report posterior samples from blockedopt SMC-ABC after a total of 32,865 and 3,337,640 model simulations. We also report the posterior samples from standard SMC-ABC after a total of 30,436 model simulations. True parameter values are indicated with black vertical lines.
5.4 Resource requirement metrics
The metrics reported in Table 1 correspond to four models and have been measured in kiloJoules (kJ) using the PowerAPI package (Bourdon et al., 2013). The experiments were run on a machine with a CPU Intel® CoreTM i7-4790 CPU @ 3.60GHz and 16GB of DRAM. Memory cost and peak memory usage results have been measured by the tracemalloc Python library. These results highlight the advantage brought by SeMPLE, which can be run on a minimal configuration while maintaining good performance. For instance, in the Two Moons case, the memory requirement (Peak Memory Usage) for SeMPLE is 13 times lower than SNPE-C and 27 times lower than SNL. The gain is even larger for the other models in Table 1. Similarly, for the Ornstein-Uhlenbeck model the energy consumption (CPU+DRAM) is more than 10 times smaller with SeMPLE.
5.5 Results summary
In Table 2 we give a summary of the main results from all experiments where an objective comparison with NN-based methods was possible, including those in Supplementary Material. Instead, for the Lotka-Volterra model we did not run NN-based methods and we reported comparisons with SMC-ABC both in section 5.3 and in the Supplementary Material. Generally SeMPLE returns accurate inference, either by producing the smallest metric value (the smaller the better), as for Two Moons and Hyperboloid, or being second best after SNL, as for Bernoulli GLM and Ornstein-Uhlenbeck, also considering the larger variation of the SNPE-C results. However, even when SNL produces a smaller Wasserstein distance, it has to be balanced with respect to the running time. For Bernoulli GLM, SNL is between 18 and 75 times slower than SeMPLE, depending on the number of rounds, as SNL has to perform an expensive neural networks training at each round. For the same reason, when looking at Bernoulli GLM, SNPE-C is between 8 and 26 times slower than SeMPLE, again depending on how many times it needs to be retrained. When considering the Ornstein-Uhlenbeck model, the Wasserstein distances have a good degree of overlap between SNL and SeMPLE, but SNL is between 9.5 and 10.5 times slower than SeMPLE, while SNPE-C is between 17 and 22 times slower than SeMPLE. In Table 2 we highlight whether an experiment has a unimodal or a multimodal target, as this determines the type of metric we report. Namely we use the C2ST metric for multimodal targets and the Wasserstein distance for unimodal targets. The reasons for this are now discussed, however notice that in the Supplementary Material we report results for both metrics in each experiment. In our experiments, we found the C2ST metric (advocated in Lueckmann et al., 2021) useful to evaluate the inference quality with multimodal targets, when draws from a reference posterior are available. However, we also found that at times C2ST produced contradicting results with unimodal targets. An example is the Binomial GLM model considered in Supplementary Section 7. There, it is obvious that the marginal posteriors returned by SeMPLE are accurate and that inference improved between the first two rounds, however C2ST showed that inference was apparently deteriorating for SeMPLE from round $r=1$ to $r=2$, which was not the case in reality. We also computed Wasserstein distances, which could confirm that the inference quality via SeMPLE did improve from round 1 to round 2. However, for multimodal targets, it is true that Wasserstein distances may not be a suitable tool, see e.g. Balaji et al. (2019). We therefore warn the reader from blindly relying on a single performance metric. It is also important to examine marginal and joint posterior densities.
Table 2: Median metric results, across 10 independent runs (brackets include minimum and maximum values), and median runtime for posterior inference. Best (lowest) values are in bold. Notice, “Metric (final)” reports the metric at the final inference round, instead “#Model sims” gives the total number of model simulations across all rounds for a single run, and similarly for “Runtime”. SNPE-C and SNL used the default SBIBM setup except when the number of algorithm iterations is reduced from 10. The C2ST metric is reported for multimodal posterior distributions and the Wasserstein distance for unimodal posterior distributions according to the discussion in Section 5.5.
Method Metric (final) #Model sims Runtime (minutes) #Rounds
Two moons (multimodal)
SNL C2ST=0.62 [0.56, 0.92] 10K 19 [17, 31] 4
SNL C2ST=0.60 [0.56, 0.68] 10K 22 [21, 23] 10
SNPE-C C2ST=0.59 [0.54, 0.66] 10K 8.1 [7.1, 12] 4
SNPE-C C2ST=0.57 [0.52, 0.60] 10K 15 [13, 17] 10
SeMPLE C2ST=0.54 [0.50, 0.58] 10K 8.3 [7.9, 8.8] 4
Hyperboloid (multimodal)
SNL C2ST=0.90 [0.84, 1.0] 40K 35 [30, 39] 4
SNL C2ST=0.90 [0.86, 0.95] 40K 50 [45, 57] 10
SNPE-C C2ST=0.76 [0.74, 0.81] 40K 100 [76, 150] 4
SNPE-C C2ST=0.74 [0.70, 0.75] 40K 124 [110, 151] 10
SeMPLE C2ST=0.58 [0.56, 0.60] 40K 14 [13, 17] 4
Bernoulli GLM (unimodal)
SNL Wasserstein=0.58 [0.44, 0.71] 10K 22 [22, 24] 2
SNL Wasserstein=0.69 [0.46, 0.89] 10K 91 [91, 92] 10
SNPE-C Wasserstein=0.94 [0.65, 1.6] 10K 9.5 [4.9, 15] 2
SNPE-C Wasserstein=0.71 [0.49, 1.78] 10K 31 [25, 40] 10
SeMPLE Wasserstein=0.77 [0.43, 1.0] 10K 1.2 [0.98, 1.2] 2
Ornstein-Uhlenbeck (unimodal)
SNL Wasserstein=0.067 [0.057, 0.23] 40K 56 [54, 60] 4
SNL Wasserstein=0.12 [0.079, 0.20] 40K 62 [58, 71] 10
SNPE-C Wasserstein=0.76 [0.27, 1.0] 40K 100 [83, 150] 4
SNPE-C Wasserstein=0.34 [0.19, 0.48] 40K 130 [100, 150] 10
SeMPLE Wasserstein=0.15 [0.11, 0.28] 40K 5.9 [3.7, 14] 4
6 Conclusion and perspectives
We have addressed the challenging question of Bayesian inference in the absence of explicit likelihoods. We have investigated mixture models through the GLLiM method, and shown that their expressivity could account for multivariate distributions, while their much simpler structure made them more amenable to efficient learning than NN solutions. The resulting new inference strategy named SeMPLE yielded inference that was on par with inference from the state-of-art neural-network (NN)-based methods that we examined (SNL and SNPE-C), while being more energy and memory efficient. For the considered examples, SeMPLE produced inference of superior or similar quality than NN-based methods, and when it did occasionally under-perform, it was by a negligible amount that can be largely be compensated by the much smaller runtime. In fact, in all tested cases, SeMPLE displayed much lower running times and memory requirements. The energy gains appear to increase with the complexity of the model simulator, and we expect to obtain even larger gains with real-life simulators. For example, for the Ornstein-Uhlenbeck model, SeMPLE achieved a 17-fold runtime speedup compared to SNPE-C. Although very preliminary, we hope this attempt, to meet the constraints of Green AI (Schwartz et al., 2020), will open the way to the development of more methods that can achieve accurate posterior inference, balancing the environmental impact of growing energy costs with the obtained results quality.
Of course, SNL and SNPE-C could certainly be improved for each specific case study, but we did not examine or fine-tune SNL and SNPE-C on the many other possible configurations, as this can be influenced by many factors. While examining several case studies implemented in SBIBM, we restricted to simulation settings provided in the SBIBM package, including reusing the same NN architectures and training setups, see Supplementary Section 1 for details.
Some open questions, that can be addressed in future research, follow. SeMPLE can benefit from the summarization of data when the dimension of $\bm{y}$ is large, just like other SBI methods. In addition to reducing the dimension, using summary statistics (as we did with the Lotka-Volterra model) can also improve the GLLiM performance. Indeed the performance of SeMPLE depends on the availability of good first approximations of the posterior and likelihood, which themselves depend on the ability of GLLiM to find probabilistic mappings between parameters and observations, in a regression-like manner using a Gaussian mixture structure. The construction of informative summary statistics is an important, but independent, research question that affects not only standard ABC methods, but also SBI methods based of neural networks, e.g. Chen et al. (2021); Radev et al. (2023b). Finally, following the discussion in Section 5.5, we wish to emphasize that measuring the quality of posterior inference in SBI is certainly an important question deserving further research, and that multiple metrics should be considered when evaluating experiments.
Acknowledgements
We thank Emeline Perthame (Institut Pasteur) for assistance with the xLLiM package. HH and UP acknowledge funding from the Swedish National Research Council (Vetenskapsrådet 2019-03924). UP acknowledges funding from the Chalmers AI Research Centre.